OpenStudy (mikezack123):

The length of a rectangle is four times its width. If the perimeter of the rectangle is 100 in, find its area.

OpenStudy (mikezack123):

@jdoe0001 @jim_thompson5910 please help :)

OpenStudy (mikezack123):

nvm got it its 400 inches squared

OpenStudy (mikezack123):

thanks

OpenStudy (anonymous):

u know the perimeter = 100 you get 4w*w = 100 4w^2=100 w^2 = 25 w = 5 area = base * hight